OpenStudy (anonymous):

cotxsec^4x=cotx+2tanx+tan^3x

OpenStudy (anonymous):

Solve by trig equations. Make the left equal the right.

OpenStudy (cwrw238):

you might try converting each term into sin's or cos's

OpenStudy (anonymous):

\[\sec ^{4}x=(1+\tan ^{2}x)^{2}\]

OpenStudy (mathmale):

Typically, the instructions for this type of problem would be, "Prove the following identity using trig identities." I would immediately convert cot x into (cos x)/(sin x) and tan x into (sin x)/cos x) for simplicity, although we don't HAVE TO do that. Here's what your equation would look like now: \[\frac{ \cos x }{ \sin x }*\frac{ 1 }{ \cos ^{4}x }=\frac{ \cos x }{ \sin x }+2\frac{ \sin x }{ \cos x }+\frac{ \sin ^{3}x }{\cos ^{3}x .}\] I'd suggest you try factoring (cos x) / (sin x) out of the right side, just as it has been factored out of the left side.

OpenStudy (anonymous):

\[\cot x(1+2\tan ^{2}x+\tan ^{4}x)\]

OpenStudy (anonymous):

\[\cot x=\frac{ 1 }{ \tan x }\]

OpenStudy (anonymous):

\[\cot x(1+2\tan ^{2}x+\tan ^{4}x)=(\cot x+2\tan ^{2}x \cot x+\cot x \tan ^{4}x)\]=\[\cot x+2 \tan x+\tan^3x\]

OpenStudy (cwrw238):

or if u write s = sin x and c = cos x ( because i;m lazy) referring to mathmales post: left hand side = c / s * c^4 = 1 / s * c^3 right hand side = c/s + 2s/c + s^3/c3 convert to single fraction gives c^4 + 2s c^2 + s^4 ------------------ s * c^3 = (c^2 + s^2)(c^2 + s^2) ------------------- = 1 / s * c*3 (as s^2 + c^2 = 1 s * c^3 so LHS = RHS